OpenStudy (anonymous):

Calculate the slant height for the cone. Round to the heart tenth. Diameter=8cm height=9cm Answers: A)10.2 B)11.4 C)9.8 D)12.0

OpenStudy (academicgurusinc):

Alright so this is relatively simple. First of all the formula for slant height is \[\sqrt{r^2+h^2}\] where "r" represents radius and "h" represents your height. In your case, you have the height so that's not an issue, but you do not have the radius. However, since you have the diameter, you can simply divide it by 2 to get our diameter if you remember. So 8/2 = 4 ... Now it's just a matter of plugging everything into the formula. So \[\sqrt{4^2+9^2}\] now you'll simplify so you get \[\sqrt{16+64}\] and when you take a square root of that, you'll get 9.8 I hope that helps. If you found this helpful, I encourage you to subscribe to our youtube channel, to stay current with all of our new videos: https://www.youtube.com/channel/UCYiI7SmkU4_vhdSzKBWsifg You may also want to check out our Math Challenge contest, for a chance to win a $50 Visa gift card: https://www.youtube.com/watch?v=2Dbu-R_Hj7E Regards, Academic Gurus Inc.

OpenStudy (academicgurusinc):

**To get our radius**